OpenStudy (anonymous):

simplify completely (12X+36)/X^2-4X-21 and find the restrictions on the variable?

OpenStudy (lalaly):

12(x+3)/(x-7)(x+3) =12/(x-7)

OpenStudy (anonymous):

= 12(x+3)/(x+3)(x-7) =12/(x-7) \[x \neq-3,x \neq7\]